OpenStudy (anonymous):

10x^2 + 23xy - 42y^2

OpenStudy (anonymous):

Differentiate?

hartnn (hartnn):

factorize ?

OpenStudy (anonymous):

factoring...sorry

OpenStudy (kira_yamato):



OpenStudy (kira_yamato):

(5x - 6y)(2x+7y)

OpenStudy (anonymous):

10x^2 + 23xy - 42y^2 (5x - 6y)(2x + 7y)

OpenStudy (anonymous):

Explanation: 10x^2 + 23xy - 42y^2 = 10x^2 + 35xy - 12xy - 42y^2 = 5x(2x + 7y) - 6y(2x + 7y) = (5x - 6y)(2x + 7y)